INAF LBC+MODS queue observing, 2011 October 19-20

Observer: K. Boutsia, M. Fumana, P. Santini
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson, O. Kuhn
Instrument: LBC+MODS

Summary

Clear skies at the beginning of the night but with very variable seeing from 0.7'' to 1.3''.
0.5h lost due to telescope problems - some actuators were off and had to be rebooted.
~1hr lost du to problems switching fromLBC to MODS - the shell ripped,the guider and wfsc had to be rebooted.
Lost 2x400sec images because the telescope jumped as we were transiting
The LBC team has reset a CCD controller manually and it has recovered automatically.
First INAF science night with MODS. A full set of OBs has been tested including calibrations. The data looks promising.

Data Summary

Project Objects Filter ExpTime Notes
Bragaglia Ber81 Bbess - Vbess / Ibess 20min the seeing was <0.8'', the OB finished at airmass 1.45 (lower than the requested limit 1.3) but we consider it within spec
Bragaglia Ber98 Bbess - Vbess / Ibess 20min the seeing was <1'' - within spec
Bragaglia Ber99 Bbess - Vbess / Ibess 20min the seeing was ~1.3 for the last blue image but all red images have <1'' - within spec
Clementini AndXXV Bbess - Vbess 1hr 20min the seeing was <1'' for most images - within spec
Clementini AndXXVII Bbess - Vbess 1hr 20min the seeing was <1'' for most images - within spec
STD Stetson_L92 Bbess - Vbess / Ibess   1 position
STD SA113 Bbess - Vbess / Ibess   4 positions
STD SA98 Bbess - Vbess / Ibess   4 positions
Donnarruma MAC0717b2 r_sdss ~1hr the seeing was <1.2'' - STD + flat +lamps + bias obtained

Details

12:51 Dome is open and moving to take flats

01:05 SkyFlats
  • BV 0 (2sec-1sec)
  • BV_180 (1sec - 1sec)
  • BI_0 (1sec-1sec)
  • BI_180 (3sec-1sec)
  • the blue counts are a bit low but in the red channel both filters should be fine

01:24 Slew to pointig star WT10_198
  • dofpia,/first
  • it seems that the telescope is not tracking in the red channel - intstead it is very bad astigmatism (??)
  • repeat dofpia after manually setting global offsets Z=-0.8
  • stopped after the 9th iteration - it is almost converging

01:45 lbcrangebal on WT10_198
  • it does not work - it gives CA=0.0 and CE=0.0
  • it was a network problem - we switched pc (obs3-> obs2) and now it is working

01:55 Slew to SA113 for standard calibration (elev~53 , airmass=1.23)
  • dither in 4 positions

02:02 before going to our target we run again lbcrangebal on WT10_198

02:06 Slew to ber81 - dofpia
  • dimm 0.9''

02:14 ber81.ob
  • after first image, red channel clearly out of focus - stop OB
  • dofpia

02:25 ber81.ob again
  • red images still look bad - elongated
  • we look at the PSF gui and some actuators are off

02:35 OB interrupted so that they can restore the actuators
  • in the meantime we do dofpia
  • blue channel started really bad but it recovered
  • run dofpia once again - actuators are fine now and both channels converge

02:46 ber81.ob again
  • half an hour lost due to telescope problems (actuators)
  • images look fine now
  • seeing measured on the images is <0.8''
  • the OB finishes at airmass~1.45 (limit requested by the PI is 1.3 but with such good seeing we will consider them within spec)

03:01 Slew to ber98.ob - dofpia

03:11 ber98.ob
  • seeing measured on the images <1.0''

03:25 Slew to Stetson_L92_1 (airmass ~ 1.5)

03:34 while we are uploading the second OB - LBC crashes (probably because the OB is not correct).
  • it automatically restarts

03:35 Stetson_L92_2/3/4/5 (airmass=1.5)

03:51 preset to ber99 - dofpia

03:58 ber99.ob
  • wind is picking up and seeing is getting worse
  • the last blue image has fwhm~1.5 but all red images are <0.9'' (PI limit is 1.3, so we consider OB within specs)

04:14 slew to M31_AndXXV_400s_1 - dofpia
  • dimm ~ 1.5''

04:20 M31_AndXXV_400s_1.ob
  • seeing measured on the images is always <1.0''

04:43 slew to M31_AndXXVII_400s_1 - dofpia
  • dimm ~ 1.4''

04:47 M31_AndXXVII_400s_1.ob
  • seeing measured on first images is <1.0''
  • last images have seeing <1.3''
  • dimm 1.4''

05:11 slew to M31_AndXXV_400s_2 - dofpia

05:16 M31_AndXXV_400s_2.ob
  • seeing measured on images: blue 1.25'' / red 1.15''
  • last images have seeing ~0.7'' (blue) and ~0.9'' (red)

05:41 slew to M31_AndXXVII_400s_2 - dofpia * dimm 1-1.2''

05:47 M31_AndXXVII_400s_2.ob
  • on of the blue images has 1/5 of the 4th chip saturated - F. Ped, will change clocking settings to see if this is the reason
  • during the last exposure the camera has jumped and all sources are double in both channels.
  • this probably happened because at 6:00UT the source crosses the meridian! the azimuth tracker has a problem(?)- the elevation though is not that high (~78)

06:12 repeat OB - M31_AndXXVII_400s_2.ob to see if the guiding problem persists
  • this time all images are good and measured seeing is <0.8''
  • dimm seeing ~0.75''

06:36 The OB has finished but before moving on F. Ped. will change the clocks to see if this is that caused the saturation on the 4th chip- blue channel (lbcb*.054738.fits)
  • TurnOFF / Turn ON LBC to implement changed clocks...

06:44 slew to M31_AndXXV_400s_3 - dofpia

06:47 M31_AndXXV_400s_3.ob
  • seeing measured on first mages: blue 0.65'' / red 0.6''
  • seeing measured on images: blue <0.75'' / red <0.72''

07:09 slew to standard Stetson_L92_1/2 (two OBs only)
  • airmass 1.21

07.22 bias - to verify that changed clocks did not introduce problems

07:24 slew to M31_AndXXVII_400s_3 - dofpia

07:30 M31_AndXXVII_400s_3.ob
  • seeing measured on first mages: blue 0.86" / red 0.84"
  • seeing measured on images: blue <0.91'' / red <0.88''

07:53 slew to M31_AndXXV_400s_4 - dofpia
  • dimm seeing 0.8

07:57 M31_AndXXV_400s_4.ob
  • seeing measured on second mages: blue 0.88" / red 0.84"

08:20 slew to M31_AndXXVII_400s_4 - dofpia * dimm seeing xxxx

08:25 M31_AndXXV_400s_4.ob
  • seeing measured on second mages: blue 0.84" / red 0.90"
  • seeing measured on images ~0.9''
  • seeing measured on images ~1''

08:48 slew to standard Stetson_L92_1 (one OB only)
  • airmass 1.48

08:55 Slew to SA98 for standard calibration (airmass=1.86)
  • dither in 4 positions

09:02 Before switching to MODS F. PEd.&Co need to do a test
  • during a 180sec exposure, the tracker has been manually disabled and it has recoverd on its own.
  • in the meantime the telescope kept guiding.
  • the test is considered successful

09:10 SWITCH TO MODS

09:30 We are switched to MODS but images do not arrive at the AGW.
  • according to an old issue track that might be caused by the AZcam.

10:16 AZcam computer has been reboot.
  • a preset has been send but the telescope sais that 'no star is found' - the star is there - we see it (active mode has failed)

10:18 repeat preset in acquire mode.
  • the computers that run the guider and the wfsc have problems - they have been rebooted.

10:36 restart GCS

10:43 we are collimating and we can start with MODS * wind is pickinh up again - seeing is getting worse (1.5'')

10:46 slewing to the target - MAC0717 (mods.7189.201130b2.mms) - Donnarumma

10:49 taking the throughslit image (through the aquisition script) - acqMODS *acq
  • MACS0717 - throughslit - aquisition (0001)
  • field image (0002)
  • to run modsAlign we NEED the mms files in one directory.
  • we cp all mms files in Donnarumma and we run the command (from the same directory)
modsAlign mods.7189.201130b2.mms mods1r.20111020.0001.fits mods1r.20111020.0002.fits - it does not work from INAF account nor from LBTO
  • we add absolute path that works on the LBTO account
modsAlign mods.7189.201130b2.mms /newdata/mods1r.20111020.0001.fits /newdata/mods1r.20111020.0002.fits - it does not work either on INAF account but it does work from LBTO
  • we go on with the mask alignment using the LBTO account
  • after centering the mask - the script exited without giving us the possibility to send the preset
  • we send the preset manually from the MODS interface: offsetpointing -0.2026 2.167 -0.952 detxy rel
  • in the throughslit all stars look reasonably aligned

11:05 execMODS *.obs
  • seeing on the MODS guider - 1.2''
  • during the first exposure the seeing si 1.1''-1.2''
  • on the first image we see only the specta of the reference stars - no obvious spectra in the other slits.

12:01 dimm 0.9''

12:16 stop during 2nd exposure (800sec missing)
  • we checked new data and it actually saved a 2250sec image (mods1r.20111020.0004.fits) and a 1381.9sec image (mods1r.20111020.0005.fits).
  • this means that if an exposure is interrupted with the STOP command a file is saved with the integration so far - you don't loose the whole image (very useful!!)

12:22 acqMODS standard_star.acq
  • we did find an appropriate guide star using LBTview
  • the preset fails because it does not manage to collimate on the star (it is becoming bright)
  • we collimate on a nearby bright star and we preset again - the preset fails
  • we go ahead and take the through slit image because it is a standard and we use the 5'' slit so we should have all flux in the slit.
  • 15sec as acquisition time is too much - tomorrow we will try 2sec
  • the star is rather centered and we go ahead with the standard_star.obs
  • the standard spectrum with 120sec has a maximum of 30000counts - this should be ok.

12:40 Closing telescope - moving in MODS calibration UNIT
  • we need to check flats exposure times and whether the lamps are the correct ones for this grating - we do have a script for flats and lamps
  • tomorrow we will need to make a cal script also for flats without the slits (for pixel to pixel corrections)
  • before executing biases we need put MODS to sleep execMODS ~MODSeng/modsScripts/Support/modsSleep.pro
  • after this we run the bias script

14:00 END OF NIGHT
  • flats and arcs for MOS mask
  • bias
  • flats and arcs for the 5'' slit were obtained on 20111021UT

-- KonstantinaBoutsia - 20 Oct 2011
Topic revision: r12 - 27 Nov 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback